# Evaluating Human-AI Collaboration: A Review and Methodological Framework

### v1

@@ 7 Limitations in Investigating Behavioral Factors
Our study provides a thorough framework to evaluate Human-AI Collaboration (HAIC) in different sectors, emphasizing goals, interaction dynamics, and task allocation. However, it does not explicitly examine behavioral aspects as a distinct main category of analysis. The exclusion of behavioral aspects of HAIC is not because they are unimportant but is due to methodological and scope concerns.
The study of behavioral aspects in the context of HAIC involves a multidisciplinary approach, incorporating knowledge from psychology, sociology, cognitive science, and information technology to address its complexity. The complexity of the research requires interdisciplinary collaboration and skills outside the current study’s context.
Behavioral aspects are important but can vary significantly and depend greatly on the context. The wide range of human behavior, shaped by cultural, psychological, and situational elements, presents a difficulty in establishing a uniform or universal evaluation framework for this topic. Extensive empirical study, including longitudinal studies and ethnographic methodologies, like those in Yang, Carroll, & Dragan (2022) [55], would provide a richer picture of how AI adaptability impacts human trust over time and how trust, in turn, shapes collaborative goals.
Our focus on the structural elements of HAIC, like goals, interaction, and task allocation, enables a more objective examination based on observable and quantifiable results. Behavioral elements demand methodological techniques and analytical frameworks different from those used in our study due to their subjective interpretations and qualitative judgments.
Behavioral variables are recognized as crucial for the success and efficacy of HAIC systems. We support more study focused on understanding these factors thoroughly. These studies could provide useful insights into the mutually beneficial connection between human views, beliefs, and behaviors and AI collaboration. This would help in creating more sophisticated, efficient, and human-centered HAIC systems.
Behavioral aspects are crucial in understanding and optimizing HAIC. However, they were intentionally excluded as a primary emphasis in this study to keep the scope focused and manageable. We urge further study to fill this gap, expanding on the groundwork established by our approach to investigate the complex interaction between human behavior and AI collaboration.
@@ 8 Conclusion
This paper has explored the current state in the evaluation of HAIC and has illuminated a notable gap in the current HAIC evaluation methodologies: the absence of a framework that effectively integrates qualitative metrics alongside quantitative ones, applicable across a wide range of sectors. Our investigation revealed that existing evaluation practices are predominantly domain-specific, lacking the structured adaptability required for comprehensive HAIC assessment. The framework proposed herein endeavors to address this void by offering a structured approach to HAIC evaluation, one that is predicated on the mode of collaboration rather than the sector, application, or AI method used. This decomposition of evaluation into smaller, manageable components presents a novel way to scrutinize HAIC in its multifaceted nature.
The significance of our framework lies in its universal applicability, affording a comparative lens to examine different collaboration methods. By transcending domain-specific limitations, it paves the way for a standardized evaluation approach that could be adopted across various sectors, enhancing the consistency and reliability of HAIC assessments.
While the framework stands as a theoretical construct, its real-world implications remain untested. The future trajectory of our research is clear: to implement this framework as a software platform and apply it across diverse sectors and collaboration modes to validate its efficacy and adaptability. Such empirical endeavors will be crucial in refining the framework and ensuring its robustness.
Our study acknowledges its limitations, particularly the omission of behavioral factors and ethical considerations from the framework. These elements, while outside our current scope, are critical dimensions of HAIC that warrant attention in future research.
As we look to the horizon, our belief is steadfast in the potential of this framework to significantly advance the methods of evaluating HAIC. As this interdisciplinary field evolves, a standardized framework becomes increasingly crucial for its adoption from industry to society. We anticipate that our contribution will serve as a catalyst for future studies, spurring the development of HAIC evaluations that are as dynamic and intelligent as the collaborations they aim to measure.
@@ Acknowledgments
The work leading to these results has received funding from the European Union’s Horizon Europe research and innovation programme under Grant Agreement No. 101120218, project HumAIne (Hybrid Human-AI Decision Support for Enhanced Human Empowerment in Dynamic Situations). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ 5 Limitations in Investigating Behavioral Factors
Our framework for evaluating Human-AI Collaboration (HAIC) emphasizes goals, interaction dynamics, and task allocation while excluding behavioral aspects due to methodological and scope constraints. Investigating behavioral factors involves interdisciplinary approach from psychology, sociology, cognitive science, which require complex and varried methodologies. While these factors are crucial for understanding trust and collaborative goals over time—exemplified by studies such as Yang, Carroll, & Dragan (2022) [49]—our study focuses on structural elements that are more readily quantifiable. We advocate for future research to explore behavioral dimensions in-depth to complement our objective framework.
@@ 6 Conclusion
This paper makes a two-fold contribution to HAIC evaluation. First, it offers a comprehensive literature review of the existing methodologies and their limitations. Second, it introduces a structured framework that integrates key evaluation factors identified in the literature and provides a standardized method for assessing HAIC across various domains.
A key gap in current methodologies is the absence of a framework that integrates both qualitative and quantitative metrics across diverse sectors. The framework proposed herein addresses this limitation by offering a structured evaluation approach based on the mode of collaboration rather than a specific sector or AI application. By breaking down evaluation into smaller, manageable components, it provides a flexible method to assess HAIC across different contexts.
Although the framework is theoretical at this stage, future research will focus on empirical validation through software implementation and real-world applications across various sectors, ensuring its adaptability and robustness. Additionally, our study acknowledges limitations, particularly the exclusion of behavioral and ethical considerations from the evaluation framework. These dimensions, though outside the current scope, are essential aspects of HAIC that should be explored in future research.
@@ Acknowledgments
The work leading to these results has received funding from the EU’s Horizon Europe research and innovation programme under GA No. 101120218, project HumAIne. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A The Unique Challenge of Evaluating Creative and Linguistic AI
Establishing comprehensive frameworks for assessing HAIC is vital, however certain applications pose distinct challenges due to their unique nature. Two prominent examples are Large Language Models (LLMs) and Generative AI in the Arts, where AI’s impact on human creativity, expression, and communication requires specialized evaluation methodologies.
While our framework provides a foundation, traditional metrics may be inadequate for these domains. Can quantitative measures fully capture artistic collaboration or AI’s influence on human language? How do we assess ethical concerns, bias, or AI-generated content’s societal impact? These challenges call for a nuanced evaluation approach. Although we don’t address these challenges in this work, we briefly discuss them here for completeness.
A.1 Large Language Models (LLMs)
LLMs have transformed text generation, translation, and dialogue. The GPT-3 [127] family, including ChatGPT and GPT-4, demonstrates their capacity for complex NLP tasks [128]. Evaluating HAIC with LLMs requires developing methodologies beyond traditional metrics [129].
Ensuring interpretability and fairness is vital. Binz & Schulz (2022)[130] show that even high-performing LLMs may use opaque reasoning patterns. Evaluation frameworks must assess how well LLMs explain their outputs and mitigate biases. Aher et al. (2022)[50] highlight the importance of testing LLM adaptability in user interactions.
LLMs augment cognition, enhance creativity, and personalize learning. Webb et al. (2022)[131] show their potential in analogical reasoning. Douglas (2023)[132] emphasizes their impact on ideation and artistic expression, while Liu et al. (2023) [133] explore their role in education. Evaluating these aspects requires metrics for creativity, knowledge retention, and learner autonomy.
Longitudinal studies provide insights into HAIC with LLMs. Yin et al. (2023)[134] suggest tracking how AI capabilities and user trust evolve over time. Dwivedi et al. (2023) [135] address societal implications, underscoring the need for ethical AI use.
Developing robust LLM evaluation frameworks is essential. Addressing interpretability, bias, and interaction quality will help create responsible, effective HAIC systems.
A.2 Generative AI in the Arts
Generative AI is reshaping artistic expression across visual art, music, and poetry. AI moves beyond tools to co-creators, expanding artistic possibilities. Notable examples include AI-generated paintings, compositions, and interactive installations. Gmeiner et al. (2022) [136] suggest leveraging team learning frameworks to enhance AI collaboration in creative fields.
Evaluating HAIC in generative AI requires adapting traditional art metrics—such as aesthetic quality, originality, and audience reception—to consider AI’s role in the creative process [137]. Verheijden & Funk (2023) [51] stress assessing AI’s impact on ideation, communication, and artistic workflows.
New methodologies must evaluate AI-assisted artistic creation holistically. He et al. (2023)[138] propose co-creativity as a joint search process, offering insights into human-AI dynamics. Ethical concerns also arise, including copyright, ownership, and authenticity. Epstein et al. (2023)[139] call for interdisciplinary research into AI’s cultural and economic impact.
As AI-assisted art evolves, evaluation models must adapt. Future research should examine real-time, personalized experiences and how AI reshapes social interactions in creative settings [140]. Documenting the artist-AI collaboration process may become as important as assessing final works.